# thanks: Corresponding author

### v1

@@ I Introduction
An additional complication for interatomic potentials of Fe- and Ni-containing alloys is the need to account for magnetism. Since Fe and Ni are strongly ferromagnetic, alloys thereof also contain significant magnetic moments [32]. Explicitly including magnetic degrees of freedom in the potential formalism is still challenging [33], especially for alloys [34]. Implicitly accounting for magnetism by fitting to spin-polarized DFT data is common practice for pure elements like Fe [35, 36, 37], but for alloys it is still rarely attempted [38, 39]. To our knowledge, no attempt has been made to (implcitly or explicitly) consider magnetism in ML potentials for FCC HEAs. Neglecting magnetism completely for AlxCrCuFeNiy alloys prohibits studies of phase stability and segregation, since without spin-polarized DFT the ground state of Fe is FCC instead of BCC. In EAM and MEAM alloy potentials [40, 41], the elemental potentials are fitted to experimental data and hence do reproduce the correct ground state without explicitly considering magnetism, but the accuracy and reliability for studying compositional phase transitions is unclear.
There has been significant interest in MD studies of radiation damage in HEAs. This is due to favorable properties in regards to radiation resistance, like lower defect concentrations, inhibited cluster growth and low swelling [42, 43]. In AlCoCrFeNi alloys there have been studies of radiation damage both experimentally [44, 45, 46, 47] and computationally [48, 49]. However, as far as the authors are aware, in AlCrCuFeNi alloys neither experimental nor computational studies have been performed.
@@ II.1 Tabulated Gaussian approximation potential
The screening function $\phi(r_{ij}/a)$ was refitted to repulsive dimer data from all-electron DFT calculations [55] for each element pair. Additionally, the screened Coulomb potential is multiplied by a cutoff function to force it to zero well below the nearest-neighbour distance of the material, to avoid interfering with the near-equilibrium interactions described by the machine-learned part. The range of the cutoff was chosen to be 1.01.01.01.02.22.$2$ to ensure smooth transition between the all-electron calculations and short-range DFT data that was part of the training data.
@@
The GAP fitting includes several hyperparamaters used for the different descriptors of the potential. The cutoff radii of the descriptors were 5.2 Å for the two-body and the EAM-like scalar density descriptor, while 4.7 Å was chosen for the three-body descriptor. The number of sparse descriptor environments from the training structures $M$ for the two-body, EAM and three-body descriptors, were 20, 30 and 500 respectively. Another set of hyperparameters are the regularisation terms $\sigma$ for the energies, forces and virials. The default values for these parameters were; 2 meV/atom, 0.1 eV/Å and 0.2 eV, respectively. For high-energy systems, such as liquid structures, dense structures, and short-range interactions, these regularisation terms were increased with a factor ten ($5\sigma$). The full input of the GAP training are available in [LINK TO REPOSITORY TO BE ADDED UPON ACCEPTANCE]. The number of grid points used in the tabulation was 1000 for both two-body and EAM descriptors and for the three-body grid $100\times 100\times 100$.
@@ II.3 Training and testing data
  • Short-range systems where an atom is randomly inserted into a FCC lattice close to other atoms, but not too close ($\sim$>∼ >1 Å) .
@@
For alloys we created data for all binary, ternary, quaternary combinations. These were created in a way that the compositional space was sparsely uniformly sampled in order to have a wide coverage of different combinations and compositions. In addition to this, some random compositions were added later with random distortions and different volumes. Furthermore, dimer and trimer data was created for each atom pair and triplet. Additionally, known ordered or intermetallic structures were added in to the training data.
For the full quinary (five-element) composition an extensive number of random compositions were created. Additionally, the compositional space of AlxCrCuFeNiy, where $x$ is 0–1 and $y$ is 1–3 was given extra focus. This range was chosen based on assessment of typical ranges of experimentally relevant compositions. For the HEA compositions several different structure types were created, including finite temperature systems, liquid structures, short-range systems, systems including vacancies and interstitials. Furthermore, during the development of the potential an iterative approach was employed, where earlier versions of the final potential were used in MD simulations and structures from these were added in to the training set. For example, some ordered HEA structures using Monte-Carlo swapping were created in this manner and added to the training data.
@@ II.4 Monte Carlo swapping MD
In order to investigate the ordering, segregation and phase stability of the HEAs, Monte Carlo swapping MD (MCMD) simulations were performed. All MD simulations in this work were run using the lammps simulation package [66]. In Monte Carlo swapping MD (MCMD) simulations, atom swaps between different chemical types are attempted at regular intervals and the probability of a swap is based on the Metropolis criterion [67]. The MCMD simulations in this work were run in the $NPT$ ensemble at 300 K for 200 ps, with atom swap attempts every 10 timesteps and 5 swap attempts between each chemical pair. We ran MCMD simulations in different compositions starting from both FCC (500 atoms) and BCC (686 atoms) structures. Additionally, an equiatomic polycrystalline alloy sample with around 42 000 atoms containing four grains was simulated. Due to the increased system size we increased the swapping attempts by a factor of 100 and ran the simulation for 20 ps.
@@ II.5 Overlapping cascade simulations
During cascade simulations, the boundaries of the system were connected to a thermostat ($NVT$) to simulate the dissipation of temperature to the surrounding material. Self-interaction over the periodic boundary was limited by halting and restarting the cascade event if an atom with over 10 eV energy crossed the simulation cell border. The cascade event was simulated for 20 ps, with an additional 10 ps of relaxation ($NPT$) back to the initial 300 K temperature and zero overall pressure. The final configuration was analysed using the Wigner-Seitz (WS) method [71] to determine the number of Frenkel pairs produced from the cascade events. Furthermore, DXA [72] analysis was performed, implemented in OVITO [73], in order to find the types and lengths of the dislocations. In order to investigate the differences in clustering we performed a cluster analysis for both interstitial and vacancy clusters, where the cutoff was chosen to be the midpoint between the third and second nearest neighbor distances ($r_{\mathrm{2.5nn}}$). Five different compositions of the AlxCrCuFeNiy HEA were considered in the overlapping cascades, which are listed in Table 1. These compositions were chosen to control the Al and Ni contents separately, thereby allowing us to investigate element-specific effects.
@@ III.1 Basic properties and validation
$B$ 148 $\pm$ 5 142
$E_{\mathrm{mix}}$128 $\pm$ 3 -90 $\pm$ 9 -66
$E_{\mathrm{form}}^{\mathrm{FCC}}$20 $\pm$ 3 20 $\pm$ 9 44
$E_{\mathrm{form}}^{\mathrm{BCC}}$5 $\pm$ 7
@@
$B$ 154 $\pm$ 5 151
$E_{\mathrm{mix}}$144 $\pm$ 3 -115 $\pm$ 7 -111
$E_{\mathrm{form}}^{\mathrm{FCC}}$54 $\pm$ 3 -25 $\pm$ 7 -20
$E_{\mathrm{form}}^{\mathrm{BCC}}$27 $\pm$ 7
@@
$B$ 160 $\pm$ 4 159
$E_{\mathrm{mix}}$150 $\pm$ 4 -123 $\pm$ 10 -115
$E_{\mathrm{form}}^{\mathrm{FCC}}$72 $\pm$ 4 -44 $\pm$ 10 -36
$E_{\mathrm{form}}^{\mathrm{BCC}}$37 $\pm$ 6
@@ III.2 Mixing energies and vacancy formation energies
Fig. 5 shows a histogram of the distribution of vacancy formation energies in the equiatomic HEA. From the figure we can see that there is a large spread in the vacancy formation energies. Additionally, there is a clear ordering with regard to the chemical species of the removed atom, with Al having the highest average and the largest spread. There are a few outliers visible in the distribution, namely one Al formation energy that is quite high and a couple of Cr values that are really low. The distributions in the other alloy compositions can be found in the Supplementary material (S3). Table 4 shows the average vacancy formation energies for all compositions and for the chemical species. We can see that the average vacancy formation energy decreases when going towards the equiatomic composition. Additionally, the order between the elements remains the same in all compositions. In AlxCoCrFeNi, DFT calculations have been used to investigate the Al dependency of vacancy formation energies [84]. While not directly comparable (due to addition of cobalt, different Al contents, and different chemical potentials used), the vacancy formation energies in this work are somewhat lower than those in AlxCoCrFeNi. However, a similar trend is reported regarding an increase in the Al content, which reduces the vacancy formation energies.
III.3 FCC/BCC phase stability
@@
Figure 6: Initial and final frames of Monte Carlo swapping simulations for the five different compositions. The smaller cells show the crystal structures identified by common neighbor analysis (green FCC and blue BCC).
@@
The stability of the FCC phase for the different alloys is studied in MCMD simulations. Fig. 6 shows the initial and final frames for the different compositions starting initially with an FCC structure. From the snapshots in Fig. 6, we see the joint segregation of Fe and Cr atoms as well as an additional segregation of Cu. Most interestingly, we observe the change of crystal structure from FCC to BCC for the equiatomic case. In this case, we see a Cr precipitate decorated with Fe and the other elements mixed somewhat uniformly (apart from some Cu segregation). However, when the simulations were performed with larger box sizes (FCC 4 000 atoms and BCC 2 000) this transition from FCC to BCC was not observed. The larger systems may be able to accommodate the strain due to nucleation of a new BCC phase easier, and a significantly longer simulation (or more swaps) might be neededto allow for the formation of large enough nucleation sites of the BCC phase to trigger a phase transition of the entire simulation cell. All systems that were initially BCC remained as BCC even after the MCMD simulations.
@@
In addition to the five specific compositions, we systematically scanned the compositional space of AlxCrCuFeNiy with the smaller system size by making 100 individual compositions with varying $x$ from 0 to 1 and $y$ from 1 to 3 and looked at the final structures. We observed only the transition from FCC to BCC in a few simulations of compositions close to equiatomic concentrations. However, not every simulation close to the equiatomic composition resulted in phase transition, and in some cases the mixed FCC and BCC phases appeared. In these cases, the BCC phase always consisted of Fe-Cr systems. This suggests that at least one mechanism of the transition from FCC to BCC is the nucleation of a Fe-Cr-rich precipitate (BCC), and if this region is large enough, it may trigger the crystal structure change. Looking at the formation energies in Table 3, we can see that the smallest difference between FCC and BCC is found for the equiatomic composition. In that case, the formation energies are given for a random alloy, and thus the ordering and formation of Fe-Cr precipitates can make the phase change possible due to the small difference in formation energies. We can also see some ordering from the simulations between Fe-Cr and sometimes for Ni-Al. Experimentally, Fe-Cr enrichment and precipitation (Ni-Al as well) have been widely reported [14, 85, 86, 12, 87]. There are even reports of Fe-Cr-rich BCC nano-precipitates in experimental literature [15, 13]. Similarly, Cu segregation has also been reported for the HEA [14, 10, 11, 85, 88, 89, 87]. For the equiatomic compositions, the experimental work mostly reports the formation of BCC or coexistence of BCC and FCC with BCC being more prevalent [89, 90, 91]. The good agreement between experimental observations and our MCMD simulation results, confirms that the present AlCrCuFeNi tabGAP potential is able to provide physical insight into the mechanisms driving phase stability, ordering and segregation in AlCrCuFeNi alloys.
@@
Figure 7: Initial and final frames of Monte Carlo swapping simulations of polycrystalline AlCrCuFeNi. The smaller boxes show the crystal structures and grain boundaries from polyhedral template matching analysis, where green color represents FCC atoms, blue represents BCC atoms, red represent HCP atoms and light-gray atoms are of type ”other” with no identified crystal structure.
@@
III.4 Radiation damage
Figure 8: Defect concentration as a function of estimated dose (NRT dpa) for the different compositions.

### v2

@@ I Introduction
An additional complication for interatomic potentials of Fe- and Ni-containing alloys is the need to account for magnetism. Since Fe and Ni are strongly ferromagnetic, alloys thereof also contain significant magnetic moments [32]. Explicitly including magnetic degrees of freedom in the potential formalism is still challenging [33], especially for alloys [34].Implicitly accounting for magnetism by fitting to spin-polarized DFT data is common practice for pure elements like Fe [35, 36, 37], but for alloys it is still rarely attempted [38, 39]. In recent years MLIPs that include explicit treatment of magnetism have started to appear, mainly for iron, fixed-composition iron-based alloys and iron oxides [40, 41, 42, 33, 34, 43]. To our knowledge, no attempt has been made to (implicitly or explicitly) consider magnetism in ML potentials for FCC HEAs. Neglecting magnetism completely for AlxCrCuFeNiy alloys prohibits studies of phase stability and segregation, since without spin-polarized DFT the ground state of Fe is FCC instead of BCC. In EAM and MEAM alloy potentials [44, 45], the elemental potentials are fitted to experimental data and hence do reproduce the correct ground state without explicitly considering magnetism, but the accuracy and reliability for studying compositional phase transitions is unclear.
There has been significant interest in MD studies of radiation damage in HEAs. This is due to favorable properties in regards to radiation resistance, like lower defect concentrations, inhibited cluster growth and low swelling [46, 47]. In AlCoCrFeNi alloys there have been studies of radiation damage both experimentally [48, 49, 50, 51] and computationally [52, 53]. However, as far as the authors are aware, in AlCrCuFeNi alloys neither experimental nor computational studies have been performed. Due to the high activation of cobalt, there is interest in finding Co-free alloy compositions for the use in radiation applications. Furthermore, most studies focus on single-phase materials, meaning the radiation response of dual-phase materials is less understood.
@@ II.1 Tabulated Gaussian approximation potential
The screening function $\phi(r_{ij}/a)$ was refitted to repulsive dimer data from all-electron DFT calculations [59] for each element pair. Additionally, the screened Coulomb potential is multiplied by a cutoff function to force it to zero well below the nearest-neighbour distance of the material, to avoid interfering with the near-equilibrium interactions described by the machine-learned part. The range of the cutoff was chosen to be 1.01.02.22.2 Å to ensure smooth transition between the all-electron calculations and short-range DFT data that was part of the training data.
@@
The GAP fitting includes several hyperparamaters used for the different descriptors of the potential. The cutoff radii of the descriptors were 5.2 Å for the two-body and the EAM-like scalar density descriptor, while 4.7 Å was chosen for the three-body descriptor. The number of sparse descriptor environments from the training structures $M$ for the two-body, EAM and three-body descriptors, were 20, 30 and 500 respectively. The sparse sampling method for all descriptors in this work used uniform sampling as implemented in the GAP framework. Another set of hyperparameters are the regularisation terms $\sigma$ for the energies, forces and virials. The default values for these parameters were; 2 meV/atom, 0.1 eV/Å and 0.2 eV, respectively. For high-energy systems, such as liquid structures, dense structures, and short-range interactions, these regularisation terms were increased with a factor ten ($5\sigma$). The full input of the GAP training are available in [LINK TO REPOSITORY TO BE ADDED UPON ACCEPTANCE]. The number of grid points used in the tabulation was 1000 for both two-body and EAM descriptors and for the three-body grid $100\times 100\times 100$.
@@ II.3 Training and testing data
  • Short-range systems where an atom is randomly inserted into a FCC lattice close to other atoms, but not too close ($\sim>$1 Å) .
@@
For alloys we created data for all binary, ternary, quaternary combinations. These were created in a way that the compositional space was sparsely uniformly sampled in order to have a wide coverage of different combinations and compositions. In addition to this, approximately 60 compositions were added later with small random distortions to the atom positions and cell shape. Furthermore, dimer and trimer data was created for each atom pair and triplet. Additionally, known ordered or intermetallic structures were added in to the training data.
For the full quinary (five-element) composition an extensive number of random compositions were created. Additionally, the compositional space of AlxCrCuFeNiy, where $x$ is 0–1 and $y$ is 1–3 was given extra focus. This range was chosen based on assessment of typical ranges of experimentally relevant compositions. For the HEA compositions several different structure types were created, including finite temperature systems, liquid structures, short-range systems, systems including vacancies and interstitials. Furthermore, during the development of the potential an iterative approach was employed, where earlier versions of the final potential were used in MD simulations and structures from these were added in to the training set. For example, around 50 ordered HEA structures using Monte-Carlo swapping were created in this manner and added to the training data.
@@ II.4 Monte Carlo swapping MD
In order to investigate the ordering, segregation and phase stability of the HEAs, Monte Carlo swapping MD (MCMD) simulations were performed. All MD simulations in this work were run using the lammps simulation package [70]. In Monte Carlo swapping MD (MCMD) simulations, atom swaps between different chemical types are attempted at regular intervals and the probability of a swap is based on the Metropolis criterion [71, 72]. The MCMD simulations in this work were run in the $NPT$ ensemble at 300 K for 200 ps, with atom swap attempts every 10 timesteps and 5 swap attempts between each chemical pair. The time step of the MCMD simulations was set to 0.002 ps. We ran MCMD simulations in different compositions starting from both FCC (500 atoms) and BCC (686 atoms) structures. Additionally, an equiatomic polycrystalline alloy sample with around 42 000 atoms containing four grains was simulated. Due to the increased system size we increased the swapping attempts by a factor of 100 and ran the simulation for 20 ps.
@@ II.5 Overlapping cascade simulations
During cascade simulations, the boundaries of the system were connected to a thermostat ($NVT$) to simulate the dissipation of temperature to the surrounding material. Self-interaction over the periodic boundary was limited by halting and restarting the cascade event if an atom with over 10 eV energy crossed the simulation cell border. The cascade event was simulated for 20 ps, with an additional 10 ps of relaxation ($NPT$) back to the initial 300 K temperature and zero overall pressure. The overlapping cascade simulations used an adaptive time step [76], to ensure energy conservation during the high-energy cascades. The final configuration was analysed using the Wigner-Seitz (WS) method [77] to determine the number of Frenkel pairs produced from the cascade events. Furthermore, DXA [78] analysis was performed, implemented in OVITO [79], in order to find the types and lengths of the dislocations. In order to investigate the differences in clustering we performed a cluster analysis for both interstitial and vacancy clusters, where the cutoff was chosen to be the midpoint between the third and second nearest neighbor distances ($r_{\mathrm{2.5nn}}$). Five different compositions of the AlxCrCuFeNiy HEA were considered in the overlapping cascades, which are listed in Table 1. These compositions were chosen to control the Al and Ni contents separately, thereby allowing us to investigate element-specific effects.
@@ III.1 Basic properties and validation
$B$ - 148 $\pm$ 5 142
$E_{\mathrm{mix}}$ -128 $\pm$ 3 -90 $\pm$ 9 -66
$E_{\mathrm{form}}^{\mathrm{FCC}}$ -20 $\pm$ 3 20 $\pm$ 9 44
$E_{\mathrm{form}}^{\mathrm{BCC}}$ -5 $\pm$ 7
@@
$B$ - 154 $\pm$ 5 151
$E_{\mathrm{mix}}$ -144 $\pm$ 3 -115 $\pm$ 7 -111
$E_{\mathrm{form}}^{\mathrm{FCC}}$ -54 $\pm$ 3 -25 $\pm$ 7 -20
$E_{\mathrm{form}}^{\mathrm{BCC}}$ -27 $\pm$ 7
@@
$B$ - 160 $\pm$ 4 159
$E_{\mathrm{mix}}$ -150 $\pm$ 4 -123 $\pm$ 10 -115
$E_{\mathrm{form}}^{\mathrm{FCC}}$ -72 $\pm$ 4 -44 $\pm$ 10 -36
$E_{\mathrm{form}}^{\mathrm{BCC}}$ -37 $\pm$ 6
@@ III.2 Mixing energies and vacancy formation energies
Fig. 5 shows a histogram of the distribution of vacancy formation energies in the equiatomic HEA. From the figure we can see that there is a large spread in the vacancy formation energies. Additionally, there is a clear ordering with regard to the chemical species of the removed atom, with Al having the highest average and the largest spread. There are a few outliers visible in the distribution, namely one Al formation energy that is quite high and a couple of Cr values that are really low. The distributions in the other alloy compositions can be found in the Supplementary material (Fig. S3). Table 4 shows the average vacancy formation energies for all compositions and for the chemical species. We can see that the average vacancy formation energy decreases when going towards the equiatomic composition. Additionally, the order between the elements remains the same in all compositions. In AlxCoCrFeNi, DFT calculations have been used to investigate the Al dependency of vacancy formation energies [90]. While not directly comparable (due to addition of cobalt, different Al contents, and different chemical potentials used), the vacancy formation energies in this work are somewhat lower than those in AlxCoCrFeNi. However, a similar trend is reported regarding an increase in the Al content, which reduces the vacancy formation energies.
III.3 Point defect migration energies
In order to investigate the differences in diffusion behavior between the compositions, nudged elastic band (NEB) [91, 92] calculations were performed to calculate the vacancy and interstitial migration barriers. For the vacancies, 1000 random configurations were calculated with a vacancy migration barrier determined with regards to exchange with each chemical species. Migration barriers in random alloys are asymmetric, meaning that for each barrier we obtain two migration energies. The convergence criterion for the NEB calculations was set to 0.05 eV/Å. Fig. 6 shows the distributions of the vacancy migration energies in each composition. We can see that the increase in Al content decreases the average migration energy slightly. A similar trend has been reported in AlxFeCoCrNi and the migration energies are similar in magnitude [51]. However, the trend is much smaller in our case. Furthermore, Al is also most often represented in the tails of the distributions, and in general the distributions are wider with increased Al content.
Figure 6: Distributions of vacancy migration energies from 1000 random configurations. The number of bins in the histogram is 25. Vertical dotted lines indicate the average migration energy of exchange with each species as well as the total average (black). The numerical value of the total average is given with the standard error of the mean.
[…]
Figure 7: Distributions of interstitial migration energies from 500 individual migration energies. The number of bins in the histogram is 100.
III.4 FCC/BCC phase stability
@@
Figure 8: Initial and final frames of Monte Carlo swapping simulations for the five different compositions at 300K. The smaller cells show the crystal structures identified by common neighbor analysis (green FCC and blue BCC).
@@
The stability of the FCC phase for the different alloys is studied in MCMD simulations. Fig. 8 shows the initial and final frames for the different compositions starting initially with an FCC structure. From the snapshots in Fig. 8, we see the joint segregation of Fe and Cr atoms as well as an additional segregation of Cu. Most interestingly, we observe the change of crystal structure from FCC to BCC for the equiatomic case. In this case, we see a Cr precipitate decorated with Fe and the other elements mixed somewhat uniformly (apart from some Cu segregation). However, when the simulations were performed with larger box sizes (FCC 4 000 atoms and BCC 2 000) this transition from FCC to BCC was not observed. The larger systems may be able to accommodate the strain due to nucleation of a new BCC phase easier, and a significantly longer simulation (or more swaps) is needed to allow for the formation of large enough nucleation sites of the BCC phase to trigger a phase transition of the entire simulation cell. Therefore, we increased the number of swap attempts for the larger systems so as to match the swap attempts per atom of the small systems and ran the simulations again. With the increased swap attempts we start seeing the formation of the BCC phase in the equiatomic composition, which is inline with what we report for the smaller systems. However, the change to BCC was only partial and there is roughly the same amount of BCC and FCC atoms in the final structures. The final frames of the larger systems can be found in the supplementary materials (Fig. S2). All systems that were initially BCC remained as BCC even after the MCMD simulations.
@@
In addition to the five specific compositions, we systematically scanned the compositional space of AlxCrCuFeNiy with the smaller system size by making 100 individual compositions with varying $x$ from 0 to 1 and $y$ from 1 to 3 and looked at the final structures. We observed only the transition from FCC to BCC in a few simulations of compositions close to equiatomic concentrations. However, not every simulation close to the equiatomic composition resulted in phase transition, and in some cases the mixed FCC and BCC phases appeared. In these cases, the BCC phase always consisted of Fe-Cr systems. This suggests that at least one mechanism of the transition from FCC to BCC is the nucleation of a Fe-Cr-rich precipitate (BCC), and if this region is large enough, it may trigger the crystal structure change. Looking at the formation energies in Table 3, we can see that the smallest difference between FCC and BCC is found for the equiatomic composition. In that case, the formation energies are given for a random alloy, and thus the ordering and formation of Fe-Cr precipitates can make the phase change possible due to the small difference in formation energies. We can also see some ordering from the simulations between Fe-Cr and sometimes for Ni-Al. Experimentally, Fe-Cr enrichment and precipitation (Ni-Al as well) have been widely reported [14, 93, 94, 12, 95]. There are even reports of Fe-Cr-rich BCC nano-precipitates in experimental literature [15, 13]. Ni-Al also has a strong tendency to order in our simulations, but this combination is more often mixed with other elements, whereas Fe-Cr formed clear precipitates. Similarly, Cu segregation has also been reported for the HEA [14, 10, 11, 93, 96, 97, 95]. For the equiatomic compositions, the experimental work mostly reports the formation of BCC or coexistence of BCC and FCC with BCC being more prevalent [97, 98, 99]. The good agreement between experimental observations and our MCMD simulation results give us confidence that the present AlCrCuFeNi tabGAP potential can provide physical insights into the mechanisms driving phase stability, ordering, and segregation in AlCrCuFeNi alloys.
@@
Figure 9: Initial and final frames of Monte Carlo swapping simulations of polycrystalline AlCrCuFeNi at 300 K. The smaller boxes show the crystal structures and grain boundaries from polyhedral template matching analysis, where green color represents FCC atoms, blue represents BCC atoms, red represent HCP atoms and light-gray atoms are of type ”other” with no identified crystal structure.
@@
III.5 Radiation damage
Figure 10: Defect concentration as a function of estimated dose (NRT dpa) for the different compositions. The defect concentration is defined based on Wigner-Seitz analysis including both vacancies and interstitials.